# NLS approximation and dark solitons for the Adlam-Allen model of cold collisionless plasmas

### v2

@@ I Introduction
One of the contexts where such classes of dispersive nonlinear PDE models emerged in these first steps was in the study of plasmas, e.g., through the classic work of washimi1966propagation , exploring the propagation of ion-acoustic traveling waves of small amplitude, via the KdV equation. However, what seems to be far less well-known was that almost a decade earlier than this classic work, another remarkable model encompassing hydrodynamic waves in plasmas was proposed by Adlam and Allen (AA) in 1958 Allen1958 and was subsequently expanded upon in further work in 1960 adlam1960 , preceding by 7 years the birth of soliton theory through the classic study of zabusky1965interaction . The physical setting of the AA model featured electrons and ions in a plasma with a magnetic field component along the $z$-direction (but only dependent along the spatial variable $x$). Assuming that considerable energy was imparted to the particles in the waves Allen1958 , this model was intended towards potential applications in fusion research, as well as in astrophysical phenomena with a notable example being the solar wind meyer-vernet2007basics .
@@ II.1 Derivation of the AA model
Under the above assumptions, the AA model is governed by Maxwell’s equations:
@@
and it noted that it is the relative velocity $U_{y}$ (along the $y$-direction) between electrons and ions that drives the relative motion.
@@
and adimensionalize the equations as follows. We measure the external magnetic field $B_{z}$ in units of $B_{0}$, the density $n$ in units of the unperturbed density $n_{0}$ [see Refs. Allen1958 ; Allen )], the velocities $U_{x},U_{y}$ in units of the characteristic Alfvén velocity $V_{A}=\sqrt{B_{0}^{2}/[4\pi\mu_{0}n_{0}(m_{i}+m_{e})]}$ (where $\mu_{0}$ is the permeability of the free space), the electric field $E_{y}$ in units of $E_{0}=V_{A}B_{0}/c$ time $t$ in units of $t_{0}=(m_{i}m_{e})^{1/2}c/(|e|B_{0})$, and the space variable $h$ in units $d=\sqrt{m_{i}m_{e}/[4\pi\mu_{0}n_{0}(m_{i}+m_{e})]}$. This way, Eqs. (12)-(16) can be reduced to the following dimensionless partial differential equations (PDEs) for the normalized magnetic field $B$ and inverse density $R=1/n$ (which is thus expected to physically satisfy $R>1$), in line with adlam1960 :
@@ III.1 Multiple scale analysis
where the subscript $\ell$ denotes the phase-multiple ($\ell=0,1,\ldots$) for a given harmonic, up to the order $n$ in the expansion in $\epsilon$. More specifically, the perturbed state depends on the fast variables $x,t$ via the fundamental carrier’s phase $kx-\omega t$ only (where $k$ and $\omega$ are the frequency and wavenumber of the carrier wave). Furthermore, the harmonic amplitudes $\textbf{S}_{n\ell}(x_{m\geq 1},t_{m\geq 1})$ are slowly varying, as they depend on the slow scales $x_{m},t_{m}$.
@@ IV.2 Numerical results
We first examine the case of a small-amplitude dark soliton. For this case, we set the formal small parameter to $\varepsilon=0.01$, while we set the free parameter $\eta_{0}$ in (46) to be $\eta_{0}=|Q/8P|$. This choice allows us to directly connect the soliton amplitude with the value of $\varepsilon$. The results of this investigation are depicted in Fig. 3. In the upper panels of Fig. 3 we see the initial state ($t=0$) of the magnetic field $B$ and the corresponding one for $t=200$. The inverse density $R$ is not depicted here for the sake of a more succinct presentation, since it does not provide any additional information about the dynamics of the system (i.e., it is fairly similar to the $B$ dynamics shown). In these panels we can see that, for this choice of $\varepsilon$, the perturbation theory provides an excellent approximation of the actual solution. The soliton propagates practically intact. There is a slight oscillation of the background of the solitary wave which is practically unseen (to the naked eye). In the lower left panel of Fig. 3 this oscillation is measured by calculating the variance of the maximum value of $B$. This is found to be of $O(\varepsilon^{2})$ as is expected on the basis of the expansion, since our approach provides $O(\varepsilon)$ approximations of the exact solution. It is worthwhile to mention that in the first time-step we observe a “jump” at the value of $B_{max}$. This is a product of the (partial) incompatibility of our initial conditions with the algebraic equation of interest. More specifically, after calculating $R^{(1)}$ by integrating (17), we solve (18) to find the compatible $B^{(1)}$. However, our initial conditions are accurate only up to $O(\varepsilon)$, thus the field $B$, in order to adjust to the precision of $R$, needs to perform this $O(\varepsilon)$ initial jump which is present in all the cases we examined in this work. In the bottom right panel of the figure a contour plot of the space-time evolution of the magnetic field $B$ is provided. There we can see that the solitary wave moves practically intact at a constant velocity which is in an excellent agreement with $v_{g}$ given by (35).
Figure 3: (Color online) The small amplitude case ($\varepsilon=0.01$). Upper panels: two snapshots of $B(x,t)$ for $t=0$ and $t=200$. Lower left panel: the variation of the maximum of $B$. Lower right panel: a contour plot of the space-time evolution of $B$, showcasing the traveling wave nature of the relevant (dark solitonic) dip in the magnetic field.
Figure 4: (Color online) The case of $\varepsilon=0.05$. Upper panels: two snapshots of $B(x,t)$ for $t=0$ and $t=200$. Lower left panel: the variation of the maximum of $B$. Lower right panel: a contour plot of the space-time evolution of the magnetic field $B$.
@@
Finally, we consider the case of a(n even) larger amplitude dark solitary wave, by taking $\varepsilon=0.1$. As we can see in Fig. 5, for this value of $\varepsilon$ the background deforms more notably acquiring a bump for $t=200$. Nevertheless, even for this value of $\varepsilon$ and for the time horizon considered, the variation of the maximum of $B$ remains of the order of ${\cal O}(\varepsilon^{2})$ which validates our approach. The gradual distortion of the solitary wave solution is also evident in the space-time evolution depicted in the bottom right panel of the figure.
@@
Figure 6: (Color online) The long term behavior of the approximated solution for the large amplitude case ($\varepsilon=0.1$). Left panel: The background is clearly distorted but the main dark-soliton is intact. Right panel: the variance of $B_{max}$ for the same time range
@@ IV.3 A modified (tailored-background) dark solitary wave
In an attempt to manually construct a more accurate dark solitary wave approximation, we numerically calculate a cnoidal wave of $O(\varepsilon)$ amplitude, but for a sufficiently large domain. This is done in order to obtain a more accurate background, adequately incorporating the nonlinear oscillations, which our dark solitary wave modulates. Then, we apply an envelope modulation of the form
@@
We consider here only the large amplitude case, i.e.,  the one with $\varepsilon=0.1$. The results of this study are shown in Fig. 7. In the lower left panel of this figure we see that, as expected, the oscillation of the background decreases significantly. Specifically, the variation of the background is half of the corresponding one in Figure 5 until $t=120$, allowing for a remarkably less distorted evolution. After this point, the solution seems to settle to a more stable background state. Moreover, in the time evolution of the soliton which is depicted in the two upper panels we see that the soliton propagates rather intact, as is also verified by the lower right contour plot.
Figure 7: (Color online) Constructed dark soliton solution for $\varepsilon=0.1$. Top panels: Two snapshots of $B(x,t)$ for $t=0$ and $t=200$. Bottom-left panel: The variation of the maximum of $B$. Bottom-right panel: A contour plot of the space-time evolution of $B$
@@ V Conclusions and Future Challenges
Naturally, these types of studies pave the way for further explorations. It would be interesting to examine whether such modulated waveforms can have any bearing in experimental settings of relevance to the AA model. More practically, however, one can also focus the current studies on the case of multiple solitary waves and their interactions, as well as to explore more systematically the stability of the nonlinear periodic patterns, as has been done in the context of the KdV and related dispersive partial differential equations bottman . One can also envision variations of the present setting to variants of the AA model such as the one where the magnetic field has two components and the electric field has one component, as, e.g., in the study of Abbas2020 . Generalizations of the present settings in higher (e.g., two) dimensions would also be of interest. Such studies are currently under consideration and will be reported in future publications.

### v3

@@ I Introduction
One of the contexts where such classes of dispersive nonlinear PDE models emerged in these first steps was in the study of plasmas, e.g., through the classic work of washimi1966propagation , exploring the propagation of ion-acoustic traveling waves of small amplitude, via the KdV equation. However, what seems to be far less well-known was that almost a decade earlier than this classic work, another remarkable model encompassing hydrodynamic waves in plasmas was proposed by Adlam and Allen (AA) in 1958 Allen1958 and was subsequently expanded upon in further work in 1960 adlam1960 , preceding by 7 years the birth of soliton theory through the classic study of zabusky1965interaction . The physical setting of the AA model featured electrons and ions in a plasma with a magnetic field component along the $z$-direction (but only dependent along the spatial variable $x$). Assuming that considerable energy was imparted to the particles in the waves Allen1958 , this model was intended towards potential applications in fusion research, as well as in astrophysical phenomena, related to wave phenomena in space and solar system physics meyer-vernet2007basics .
@@ II.1 Derivation of the AA model
To derive the AA model we start from Maxwell’s equations:
@@
and it is noted that it is the relative velocity $U_{y}$ (along the $y$-direction) between electrons and ions that drives the relative motion.
@@
and adimensionalize the equations as follows. We measure the external magnetic field $B_{z}$ in units of $B_{0}$, the density $n$ in units of the unperturbed density $n_{0}$ [see Refs. Allen1958 ; Allen )], the velocities $U_{x},U_{y}$ in units of the characteristic Alfvén velocity $V_{A}=\sqrt{B_{0}^{2}/[4\pi n_{0}(m_{i}+m_{e})]}$, the electric field $E_{y}$ in units of $E_{0}=V_{A}B_{0}/c$ time $t$ in units of $t_{0}=(m_{i}m_{e})^{1/2}c/(|e|B_{0})$, and the space variable $h$ in units $d=\sqrt{c^{2}m_{i}m_{e}/[4\pi n_{0}e^{2}(m_{i}+m_{e})]}$. This way, Eqs. (12)-(16) can be reduced to the following dimensionless partial differential equations (PDEs) for the normalized magnetic field $B$ and inverse density $R=1/n$ (which is thus expected to physically satisfy $R>1$), in line with adlam1960 :
@@ III.1 Multiple scale analysis
where the subscript $\ell$ denotes the phase-multiple for a given harmonic, up to the order $n$ in the expansion in $\epsilon$. More specifically, the perturbed state depends on the fast variables $x,t$ via the fundamental carrier’s phase $kx-\omega t$ only (where $k$ and $\omega$ are the frequency and wavenumber of the carrier wave). Furthermore, the harmonic amplitudes $\textbf{S}_{n\ell}(x_{m\geq 1},t_{m\geq 1})$ are slowly varying, as they depend on the slow scales $x_{m},t_{m}$.
@@ IV.2 Numerical results
The initial conditions for the numerical study of this section are taken by inserting (46) into the perturbation-theoretic expressions of Eqs. (44)-(45) evaluated at $t=0$. We first consider the case of a small-amplitude dark soliton. Since the solution amplitude is analogous to $\varepsilon$, we set the formal small parameter to $\varepsilon=0.01$, while we set the free parameter $\eta_{0}$ in (46) to be $\eta_{0}=|Q/8P|$ (so that the factor $\sqrt{8\eta_{0}|P/Q|}=1$). This choice allows us to directly connect the soliton amplitude with the value of $\varepsilon$. The results of this investigation are depicted in Fig. 3. In the upper panels of Fig. 3 we see the initial state ($t=0$) of the magnetic field $B$ and the corresponding one for $t=200$. The inverse density $R$ is not depicted here for the sake of a more succinct presentation, since it does not provide any additional information about the dynamics of the system (i.e., it is fairly similar to the $B$ dynamics shown). In these panels we can see that, for this choice of $\varepsilon$, the perturbation theory provides an excellent approximation of the actual solution. The soliton propagates practically intact. There is a slight oscillation of the background of the solitary wave which is practically unseen (to the naked eye). In the lower left panel of Fig. 3 this oscillation is measured by calculating the temporal variation of the maximum value of $B$. This is found to be of $O(\varepsilon^{2})$ as is expected on the basis of the expansion, since our approach provides $O(\varepsilon)$ approximations of the exact solution. It is worthwhile to mention that in the first time-step we observe a “jump” at the value of $B_{max}$. This is a product of the (partial) incompatibility of our initial conditions with the algebraic equation of interest. More specifically, after calculating $R^{(1)}$ by integrating (17), we solve (18) to find the compatible $B^{(1)}$. However, our initial conditions are accurate only up to $O(\varepsilon)$, thus the field $B$, in order to adjust to the precision of $R$, needs to perform this $O(\varepsilon)$ initial jump which is present in all the cases we examined in this work. In the bottom right panel of the figure a contour plot of the space-time evolution of the magnetic field $B$ is provided. There we can see that the solitary wave moves practically intact at a constant velocity which is in an excellent agreement with $v_{g}$ given by (35).
Figure 3: (Color online) The small amplitude case ($\varepsilon=0.01$). Upper panels: two snapshots of $B(x,t)$ for $t=0$ and $t=200$. Lower left panel: the temporal variation of the maximum of $B$. Lower right panel: a contour plot of the space-time evolution of $B$, showcasing the traveling wave nature of the relevant (dark solitonic) dip in the magnetic field.
Figure 4: (Color online) The case of $\varepsilon=0.05$. Upper panels: two snapshots of $B(x,t)$ for $t=0$ and $t=200$. Lower left panel: the temporal variation of the maximum of $B$. Lower right panel: a contour plot of the space-time evolution of the magnetic field $B$.
@@
Finally, we consider the case of a(n even) larger amplitude dark solitary wave, by taking $\varepsilon=0.1$. As we can see in Fig. 5, for this value of $\varepsilon$ the background deforms more notably acquiring a bump for $t=200$. Nevertheless, even for this value of $\varepsilon$ and for the time horizon considered, the temporal variation of the maximum of $B$ remains of the order of ${\cal O}(\varepsilon^{2})$ which validates our approach. The gradual distortion of the solitary wave solution is also evident in the space-time evolution depicted in the bottom right panel of the figure.
@@
Figure 6: (Color online) The long term behavior of the approximated solution for the large amplitude case ($\varepsilon=0.1$). Left panel: The background is clearly distorted but the main dark-soliton is intact. Right panel: the temporal variation of $B_{max}$ for the same time range
@@ IV.3 A modified (tailored-background) dark solitary wave
In an attempt to manually construct a more accurate dark solitary wave approximation, we adopt the following procedure. We numerically calculate a cnoidal wave of $O(\varepsilon)$ amplitude; this periodic structure is then used as a more accurate background, adequately incorporating the nonlinear oscillations, which our dark solitary wave modulates. This is achieved by considering the initial condition $w(0)=w_{min}-\varepsilon,w_{\xi}(0)=0$ and numerically integrating the dynamical system (23) for a large enough domain of $\xi$. After acquiring the background of the wave structure, we apply an envelope modulation of the form
@@
We consider here only the large amplitude case, i.e.,  the one with $\varepsilon=0.1$. The results of this study are shown in Fig. 7. In the lower left panel of this figure we see that, as expected, the oscillation of the background decreases significantly. Specifically, the temporal variation of the background is half of the corresponding one in Figure 5 until $t=120$, allowing for a remarkably less distorted evolution. After this point, the solution seems to settle to a more stable background state. Moreover, in the time evolution of the soliton which is depicted in the two upper panels we see that the soliton propagates rather intact, as is also verified by the lower right contour plot.
Figure 7: (Color online) Constructed dark soliton solution for $\varepsilon=0.1$. Top panels: Two snapshots of $B(x,t)$ for $t=0$ and $t=200$. Bottom-left panel: The temporal variation of the maximum of $B$. Bottom-right panel: A contour plot of the space-time evolution of $B$
@@ V Conclusions and Future Challenges
Naturally, these types of studies pave the way for further explorations. It would be interesting to examine whether such modulated waveforms can have any bearing in experimental settings of relevance to the AA model. More practically, however, one can also focus the current studies on the case of multiple solitary waves and their interactions, as well as to explore more systematically the stability of the nonlinear periodic patterns, as has been done in the context of the KdV and related dispersive partial differential equations bottman . One can also envision variations of the present setting to variants of the AA model such as the one where the magnetic field has two components and the electric field has one component, as, e.g., in the study of Abbas2020 . Generalizations of the present settings in higher (e.g., two) dimensions would also be of interest. Moreover, we note in passing that experimental observation of Alfvén and ion-acoustic waves (among others) is becoming experimentally more accessible as confirmed by data collected, e.g., by the Radio Receiver Instrument of the Swarm-E satellite Bernhardt2023 . As such, associated observations of plasma waves generated, e.g., by charged space objects expand upon the need for nonlinear models and their characterization/description of such nonlinear wave propagation. Such studies are currently under consideration and will be reported in future publications.